OpenStudy (anonymous):

The body mass index , I can be used to determine on individuals risk for heart disease. An index less than 25 indicates a low risk .The body mass index is given by the formula or model I =700 w/h2 where w = weight in pounds and h is height in inches . Francis weighs 200 pounds and stands 72 inches tall what is his approximate body mass index? find an inequality describing all weights w that Francis can have and be in the low risk category Francis approximate body mass index is ? round to the nearest whole number find an inquality that describing all weights w that francis can have and be in th

OpenStudy (anonymous):

Is the expression for BMI is \[I = 700\times\frac{w}{h^2}\]

OpenStudy (anonymous):

low weight category?\

OpenStudy (anonymous):

bmi for francis = 700 * 200 / 72^2 = 27

OpenStudy (anonymous):

ya

OpenStudy (anonymous):

25 > 700w / 72^2 700w < 25*72^2 w < 25*72^2 / 700 w < 185

OpenStudy (anonymous):

For Low risk category \[700\times\frac{w}{72^2}<25\] \[w<\frac{25\times72^2}{700}\] \[w<185.14\]

OpenStudy (anonymous):

yep